QUESTION :The addition of HF provides a direct way to measure the aluminium content of various types of clay. A 4.27 gram sample of kaolinite clay, Al2(OH)4Si2O5 was treated with excess HF and produced 2.54 grams of solid precipitate of AlF3 according to the equation given below.

Al2(OH)4Si2O5 (s) + 14 HF → 9 H2O + 2 SiF4 (g) + 2 AlF3 (s)

i) List two criteria for gravimetric analysis (2)

ii) Calculate the % Al in the sample (3)
Answer #1

Answer:

i) The main two criteria behind the gravimetric anaysis

a.) The compound formed must be of high purity and low solubility with the known composition.

b.) We must be able to separate the component that precipitate in the gravimetric analysis.

ii) The calculated molar mass of Al2(OH)4Si2O5 is 258.1604 g/mol.

Hence, 258.1604 g of Al2(OH)4Si2O5 contains 53.962 g of Al.

Hence, amount of Al in the sample = (53.962/258.1604) x 4.27 g = 0.8925 g

% Al in sample = (0.8925/4.27) x 100 = 20.90 %

Thanks